In an experiment on photoelectric effect, light of wavelength 400 nm is incident on a cesium plate at the rate of 5.0 W. The potential of the collector plate is made sufficiently positive with respect to the emitter so that the current reaches its saturation value. Assuming that on the average one out of every 10^6 photons is able to eject a photoelectron, find the photocurrent in the circuit.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. λ = 400 nm, P = 5 w E of 1 photon = hc/λ = (1242/400) ev No.of electrons = 5/Energy of 1 photon = 5 * 400/1.6 * 10^-19 * 1242 No.of electrons = 1 per 10^6 photon. No.of photoelectrons emitted = 5 * 400/1.6 * 1242 * 10^-19 * 10^6 Photo electric current = 5 * 400/1.6 * 1242 * 10^6 * 10^-19 * 1.6 * 10^-19 = 1.6 * 10^-6 A = 1.6 μA.